Calculate the density of a substance weighing 77 lbs that occupies 125 cubic inches. Express in terms of kg/L^3
1
Expert's answer
2021-09-09T03:09:01-0400

1lbs = 0.454 kg

= 77 × 0.454 = 34.96kg

1L = 61.024 cubic inches

125/61.024 = 2.05L

= 34.96/2.05 = 17.1Kg/L3